# Convergence Acceleration of Favre-Averaged Non-Linear Harmonic Method

### v1

@@ Results
where $R_{z,i},i\in[1:N_{h}],$ is the RMS of the non-linear residual for the $i^{\text{th}}$ harmonic. In the following text, the residual of the mean flow and the RMS of the residuals for all harmonics, namely $R_{z}$, is used to compare the performance of the explicit and implicit RK scheme.
For the performance of the explicit scheme with multigrid, Fig. 9 shows the convergence history of FNLH for the case shown in Fig. 6 and the explicit RK scheme with V-cycle multigrid is used. It can be seen from Fig. 9 that only one multigrid level yields poor convergence. As the number of MG levels increases, the convergence rate is significantly improved, but the benefit of adding more MG levels is also diminishing. The non-linear mean flow seems to benefit more from multigrid compared to the time-linearized harmonic flow. The oscillations of residuals between 250 - 500 iterations are likely caused by the transonic flows in the rotors. As the residuals are further reduced, and the flow becomes more established; such an oscillation is no longer present.
Figure 9: Convergence history of explicit schemes with different levels of MG for the 1.5 stage compressor case.
It should be noted that as the number of MG levels increases, the computational cost of each nonlinear iteration also increases. Figure 10 shows the wall-clock time against the nonlinear residual. The wall clock is expressed in TauBench Work Units (WU). One WU is the wall clock time of the following command:
mpirun -np 1 ./TauBench -n 250000 -s 10
The computational cost of the CFD simulation in terms of WU is then expressed as $\frac{N_{\text{proc}}T_{\text{FNLH}}}{T_{1}}$. $N_{\text{proc}}$ is the number of CPU cores that are used and $T_{\text{FNLH}}$ is the wall clock time of the FNLH simulation. $T_{1}$ is the wall clock time of the aforementioned TauBench command. For the explicit scheme, although each nonlinear iteration is more expensive, 4 level of MG is still the most computationally efficient one, and using only one level of MG is not computationally efficient and also shows poor convergence. In the following text, the explicit scheme with 4 levels of MG will be used as the default explicit scheme and compared with the implicit RK schemes.
@@
Figure 11 shows the comparison of the convergence histories of 4 levels of MG and implicit RK schemes with different CFL numbers. It can be seen that in terms of non-linear iterations, implicit RK schemes show significant improvement over the explicit scheme with 4 levels of MG. In addition, the oscillations of the residuals observed between 200 and 500 iterations are not present in the implicit scheme. This is because the implicit scheme uses a full Jacobian matrix with first-order spatial accuracy for the preconditioner matrix shown in Equation 9 and Equation 10 while the explicit scheme only uses the diagonal block of this matrix. Therefore, the solution update could be in a more accurate direction for the implicit scheme to reduce the residual. When there are strong non-linearity in the flow, this is beneficial to improve the convergence and also the robustness of the solution.
Figure 11: Convergence history of implicit schemes with different CFL numbers for the 1.5 stage compressor case
An approximate Jacobian matrix is used in the current implicit scheme for the mean flow (Equation 9) and the linearized flow (Equation 10). The CFL number acts as a relaxation factor. A smaller CFL number leads to better diagonal dominance, but this will also lead to slower convergence and vice versa. Therefore, a judicious choice of the CFL number will strike a good balance in terms of convergence speed, stability, and computational efficiency. Figure 12 shows that as the CFL number is gradually increased from 10 to 50, the convergence rate improves, but the benefit of increasing the CFL also decreases. A CFL number of 40 or 50 seems to be a good choice, and this is consistent with the previous experience of developing the implicit scheme for the non-linear mean flow [25].
Figure 12: Computational efficiency of implicit schemes for the 1.5 stage compressor case.
@@
Figure 13: Impact of CPU cores on the implicit scheme for the compressor case.
With respect to computational efficiency, which can be measured in terms of the work unit, Fig. 13 shows that adding more CPUs leads to a less efficient implicit scheme, but it has marginal effect on the number of non-linear iterations. This is because the accuracy of the Jacobian matrix decreases as more CPU cores are used, and this will lead to a greater number of linear iterations to reach the convergence threshold of the linear residual. Hence, each nonlinear iteration becomes more expensive, but the resulting total number of nonlinear iterations can still be similar due to the fixed linear residual threshold.
@@
The second case is based on the MT1 turbine stage [36] and was modified to a 1.5 stage turbine in our previous work [4] to validate FNLH for hot streak migrations. The original blade count of the MT1 stage is 32 (NGV):60 (R1) and the blade counts of the redesigned 1.5-stage turbine are scaled to 30 (NGV): 60 (R1): 30 (S2). A stream tube of the geometry is simulated, and the stream tube accounts for roughly $5\%$ of the blade span. The NGV is modeled as a pair to prescribe a simplified hot streak at the NGV inlet. One layer of hexahedral elements is used to discretize the computational domain, and there are 72328 hexahedral elements in total. The average $y^{+}$ on the blade surface is around 5. The hot streak, which spans two NGV passages, has a sinusoidal shape with an amplitude of $\Delta T_{hs}$. The resulting inlet total temperature field $T_{\text{in}}$ is represented as the following:
@@
in which $T_{0}$ is the mean total temperature at the inlet and its value is 444 K, and $N_{\text{hs}}=15$, which means the hot streak will span two NGV passages. The measured peak-to-mean and minimum-to-mean temperature ratios are approximately 1.08 and 0.93 in the MT1 stage [36], respectively. In this study $\Delta T_{hs}=50$K is used. The ideal-gas model is used in this study, although in the real scenario, the working fluid in the turbine should be a mixture of air and the productions of the combustion process. The ideal gas model can still be adequate for the preliminary design and for validating computational methods. For the URANS simulation, single-passage meshes are replicated in the circumferential direction to form a $12^{\circ}$ sector. 200 time steps are used for this $12^{\circ}$ sector simulation, which corresponds to 100 time steps for the stator to sweep one passage of the rotor.
The setup for the harmonics in FNLH is summarized in Table 4 and this setup is based on our previous study in this case [4]. 16 harmonics are computed in total. To model the hot streak in S2, 7 harmonics are used. 2 fundamental modes of the hot streak and their related scattered modes are computed. Their temporal frequency is related to the shaft speed $\Omega$ and its wave number is related to the linear combination of the wave number of the hot streak and the blade count of R1.
@@
NGV $-i\text{N}_{R1}\Omega,i\in[1:16]$ $i\text{N}_{R1},i\in[1:16]$ exit, R1 potential field
R1 $i\text{N}_{NGV}\Omega,i\in[1:14]$ $i\text{N}_{NGV},i\in[1:14]$ inlet, NGV wake + hot streak
$i\text{N}_{S2}\Omega,i\in[1:2]$ $i\text{N}_{S2},i\in[1:2]$ exit, S2 potential field
S2 $i\text{N}_{R1}\Omega,i\in[1:9]$ $i\text{N}_{R1},i\in[1:9]$ inlet, R1 wake
@@
Figure 14: Convergence history of explicit and implicit scheme for the turbine case.
The explicit scheme with 4 level of MG and the implicit scheme with CFL$=50$ are compared. According to the previous test case, these two are the most efficient explicit and implicit schemes shown, respectively. The residuals of the explicit and implicit schemes are shown in Figure 14 and 30 CPU cores are used. Similarly to the previous case, the implicit scheme shows a significant speed-up of roughly 10x compared to the explicit scheme in terms of computational time. For the explicit scheme with MG, the convergence rate is degraded as the residual decreases. This indicates that the explicit scheme with MG is not effective in damping the short-wavelength errors in the computation. However, the implicit scheme shows is less affected by this issue.
In terms of parallel performance of the implicit scheme, Fig. 15 shows the convergence history with 5, 10, 20, and 30 CPU cores. For the case with 30 CPU cores, the number of elements per CPU core is roughly 2410. It can be seen that in terms of non-linear iterations, the number of CPU cores has a negligible effect. With respect to WU, it can be seen that the implicit solver becomes slightly less efficient as the number of CPUs increases. This observation is consistent with the previous compressor case.
Figure 15: Impact of CPU cores on the implicit scheme for the turbine case.
Figure 16 compares the predicted instantaneous temperature field between FNLH and URANS. For the FNLH, the implicit scheme is used. There is good agreement between URANS and FNLH on the predicted temperature field, and the hot streak is correctly transported downstream the turbine bladerows. For a quantitative comparison, Fig. 17 shows the circumferential distribution of instantaneous temperature at two streamwise locations. The first is the midway between the R1 inlet and the R1 LE, and the second is the midway between the S2 inlet and the S2 LE. It can be seen that there is good agreement between the implicit FNLH and URANS. More validations of FNLH in this case can be found in previous work [4].
@@
Table 5 compares the memory consumption of the explicit and implicit RK schemes for the test cases. It can be seen that the memory consumption of the implicit solver is roughly half of the explicit solver using 4 MG levels. Therefore, the implicit RK scheme not only delivers faster and faster convergence compared to the explicit scheme but also has significantly less memory consumption. The memory consumption of the corresponding whole annulus URANS simulations is also summarized in Table 5. The memory consumption of FNLH is roughly two orders of magnitude smaller compared to whole annulus URANS.
@@
Compressor case 3.9 2.1 309.6
Turbine case 7.4 3.2 378.1
Figure 18 shows the convergence history of the implicit scheme (CFL = 50) for the test cases, but the x-axis of the plots is the normalized time, which is defined as the wall clock time of the FNLH simulation divided by the wall clock time of execution of the full annulus URANS simulation for one revolution. It is typical to run 4-5 revolutions for URANS before collecting data from URANS. Figure 18 shows that the implicit FNLH is at least two orders of magnitude more efficient than the corresponding full-annulus URANS simulations in both test cases.
@@
Conclusions
A unified formulation has been developed and implemented in the industrial CFD HYDRA to accelerate the convergence of the FNLH method. A comparative study of the resulting explicit and implicit schemes has been undertaken. The proposed method is memory efficient as memory consumption does not depend on the number of harmonics to compute. For the implicit scheme, complex algebra is used for the matrix factorization for the purpose of re-using the look-up table for matrix factorization of the mean flow. This significantly reduces the effort of implementing the implicit FNLH scheme in parallel.
The implicit scheme is found to consistently show a speed-up of 7-10x compared to the explicit schemes with multigrid. Furthermore, the memory consumption of the implicit scheme consumes approximately $50\%$ of that of the explicit scheme with 4 levels of multigrid. The parallel performance of the implicit scheme is also studied, and it shows that the implicit solver is slightly less efficient as more CPU cores are used. However, in a typical industrial simulation, the number of elements per CPU core is at least $O(10^{4})$, therefore, the parallel performance of the implicit scheme should not be a major concern.

### v2

@@ Results
where $R_{z,i},i\in[1:N_{h}],$ is the RMS of the non-linear residual for the $i^{\text{th}}$ harmonic. In the following, the residual of the mean flow and the RMS of the residuals for all harmonics, namely $R_{z}$, is used to compare the performance of the explicit and implicit RK scheme.
For the performance of the explicit scheme with different levels of MG, Fig. 8 shows the convergence history of FNLH for this compressor case. The explicit RK scheme with the V-cycle multigrid is used, and only the residual of the finest grid level is shown. It can be seen that using only one multigrid level yields poor convergence. As the number of MG levels increases, the convergence rate is significantly improved, but the benefit of adding more MG levels is also decreasing. The non-linear mean flow seems to benefit more from multigrid compared to the time-linearized harmonic flow. The oscillations of residuals between 250 - 500 iterations are likely caused by the transonic flows in the rotors. As the residuals are further reduced and the flow becomes more established; such an oscillation is no longer present.
Figure 8: Convergence history of explicit schemes with different levels of MG for the 1.5-stage compressor case.
It should be noted that as the number of MG levels increases, each nonlinear iteration becomes more expensive. It is also important to examine the residual history against wall-clock time. Because the performance of a CFD code will depend on the computer system, one way to overcome this problem is to employ a benchmark such as the TauBench code to nondimensionalize the computational cost (see Wang et al. [38]) and express the wall-clock time in terms of the Work Unit (WU), so that the running time of different CFD codes that run on different computers could be compared. In this work, WU is computed as $\frac{N_{\text{proc}}T_{\text{FNLH}}}{T_{1}}$. $N_{\text{proc}}$ is the number of CPU cores that are used and $T_{\text{FNLH}}$ is the wall clock time of the FNLH simulation. $T_{1}$ is the wall clock time for executing the TauBench code in the local system with the following parameters: “mpirun -np 1 ./TauBench -n 250000 -s 10”, and these parameters are widely used in the CFD community.
Figure 9 shows the WU against the nonlinear residual. For the explicit scheme, although each nonlinear iteration becomes expensive as the number of MG levels increases, 4 levels of MG is still the most computationally efficient one, and using only one level of MG is not computationally efficient and also shows poor convergence. In the following, the explicit scheme with 4 levels of MG will be used as the default explicit scheme and compared with the implicit RK schemes.
@@
Figure 10 shows the comparison of the convergence histories of 4 levels of MG and implicit RK schemes with different CFL numbers. It can be seen that, in terms of non-linear iterations, implicit RK schemes show significant improvement over the explicit scheme with 4 levels of MG. In addition, the oscillations of the residuals observed between 200 and 500 iterations are not present in the implicit scheme. This is because the implicit scheme uses a full Jacobian matrix with first-order spatial accuracy for the preconditioner matrix shown in Equation 8 and Equation 13 while the explicit scheme only uses the diagonal block of this matrix. Therefore, the solution update could be in a more accurate direction for the implicit scheme to reduce the residual. When there is strong non-linearity in the flow, this is beneficial to improve the convergence and also the robustness of the solution.
Figure 10: Convergence history of implicit schemes with different CFL numbers for the 1.5-stage compressor case
In the implicit scheme, the CFL number acts as a relaxation factor. A smaller CFL number leads to better diagonal dominance, but this will also lead to slower convergence. Therefore, a judicious choice of the CFL number will strike a good balance in terms of convergence speed, stability, and computational efficiency. Figures 10 and 11 show that as the CFL number gradually increases from 10 to 100, the convergence rate improves, but the benefit of increasing the CFL number also decreases rapidly. In addition, increasing the CFL reduces the diagonal dominance and could lead to potential robustness issues for complex flows in practice. From our numerical experiment, a CFL number of 40 or 50 is found to be a good choice, and this is consistent with the previous experience of developing the implicit scheme for the non-linear mean flow [25].
Figure 11: Computational efficiency of implicit schemes for the 1.5-stage compressor case.
@@
Figure 12: Impact of CPU cores on the performance of the implicit scheme for the compressor case.
With respect to computational efficiency, which can be measured in terms of the work unit, Fig. 12 shows that adding more CPUs leads to a less efficient implicit scheme, but it has marginal effect on the number of nonlinear iterations. This is because the accuracy of the Jacobian matrix decreases as more CPU cores are used, and this will lead to a greater number of linear iterations to reach the convergence threshold of the linear residual. Hence, each nonlinear iteration becomes more expensive, but the resulting total number of nonlinear iterations can still be similar due to the fixed linear residual threshold.
@@
The second case is based on the MT1 turbine stage [39] and was modified to a 1.5-stage turbine in our previous work [4] to validate FNLH for hot streak migrations. The original blade count of the MT1 stage is 32 (NGV):60 (R1) and the blade counts of the redesigned 1.5-stage turbine are scaled to 30 (NGV): 60 (R1): 30 (S2). A stream tube of the geometry is simulated, and the stream tube accounts for roughly $5\%$ of the blade span. The NGV is modeled as a pair to prescribe a simplified hot streak at the NGV inlet. One layer of hexahedral elements is used to discretize the computational domain, and there are 72328 hexahedral elements in total. The average $y^{+}$ on the blade surface is around 5. The hot streak, which spans two NGV passages, has a sinusoidal shape with an amplitude of $\Delta T_{hs}$. The resulting inlet total temperature field $T_{\text{in}}$ is represented as the following:
@@
in which $T_{0}$ is the mean total temperature at the inlet and its value is 444 K, and $N_{\text{hs}}=15$, which means the hot streak will span two NGV passages. The measured peak-to-mean and minimum-to-mean temperature ratios are approximately 1.08 and 0.93 in the MT1 stage [39], respectively. In this study $\Delta T_{hs}=50$K is used. In this study, the ideal gas model is used. For the URANS simulation, single-passage meshes are replicated in the circumferential direction to form a $12^{\circ}$ sector. 200 time steps are used for this $12^{\circ}$ sector simulation, which corresponds to 100 time steps for the stator to sweep one passage of the rotor.
The setup for the harmonics in FNLH is summarized in Table 4 and this setup is based on our previous study [4]. 16 harmonics are computed in total. To model the hot streak in S2, 7 harmonics are used. Two fundamental modes of the hot streak and their related scattered modes are computed. Their temporal frequency is related to the shaft speed $\Omega$ and its wave number is related to the linear combination of the wave number of the hot streak and the blade count of R1.
@@
NGV $-i\text{N}_{R1}\Omega,i\in\{1,2,...16\}$ $i\text{N}_{R1},i\in\{1,2,...16\}$ exit, R1 potential field
R1 $i\text{N}_{NGV}\Omega,i\in\{1,2,...14\}$ $i\text{N}_{NGV},i\in\{1,2,...14\}$ inlet, NGV wake + hot streak
$i\text{N}_{S2}\Omega,i\in\{1,2\}$ $i\text{N}_{S2},i\in\{1,2\}$ exit, S2 potential field
S2 $i\text{N}_{R1}\Omega,i\in\{1,2,...9\}$ $i\text{N}_{R1},i\in\{1,2,...9\}$ inlet, R1 wake
@@
Figure 13: Convergence history of explicit and implicit schemes for the turbine case.
The explicit scheme with 4 levels of MG and the implicit schemes with CFL$=50$ and CFL$=100$ are compared. The residuals of the explicit and implicit schemes are shown in Figure 13 and 30 CPU cores are used. Similarly to the previous case, the implicit scheme shows a significant speed-up of roughly 10 compared to the explicit scheme in terms of computational time. For the explicit scheme with MG, the convergence rate is degraded as the residual decreases. This indicates that the explicit scheme with MG is not effective in damping the short-wavelength errors in the computation. However, the implicit scheme shows that it is less affected by this issue. Furthermore, for the implicit scheme, increasing the CFL from 50 to 100 only slightly improves the convergence rate. This is similar to the observation from the compressor case.
In terms of parallel performance of the implicit scheme, Fig. 14 shows the convergence history with 5, 10, 20, and 30 CPU cores. For the case with 30 CPU cores, the number of elements per CPU core is roughly 2410. It can be seen that, in terms of non-linear iterations, the number of CPU cores has a negligible effect. With respect to WU, it can be seen that the implicit solver becomes slightly less efficient as the number of CPUs increases. This observation is consistent with the previous compressor case.
Figure 14: Impact of CPU cores on the performance of the implicit scheme for the turbine case.
Figure 15 compares the predicted instantaneous temperature field between FNLH and URANS. For the FNLH, the implicit scheme is used. There is good agreement between URANS and FNLH on the predicted temperature field, and the hot streak is correctly transported downstream the turbine bladerows. For a quantitative comparison, Fig. 16 shows the circumferential distribution of the instantaneous temperature at two streamwise locations. The first is the midway between the R1 inlet and the R1 LE, and the second is the midway between the S2 inlet and the S2 LE. It can be seen that there is good agreement between the implicit FNLH and URANS. More validations of FNLH in this case can be found in previous work [4].
@@
Low Pressure Compression System of A Turbofan Engine
The third test case is the low pressure compression (LPC) system of a high bypass ratio turbofan engine. It consists of the fan, engine section strut (ESS), and the outlet guide vane (OGV). The computational domain of this case is shown in Fig. 17. The exit boundary of the fan blade is split into two parts, one is connected to ESS and the other one is linked to OGV. The flow in the fan passage features the passage shock towards the casing and strong three-dimensional secondary flows. The fan blade is discretized with 110 layers of hexahedral in the spanwise direction and has 2077236 hexahedral cells in total. The FNLH setup for this case is summarized in Table 5. 8 harmonics are computed in total. In particular, for the fan blade, 4 harmonics are used to represent the potential field from ESS and OGV, respectively. For ESS and OGV, 8 harmonics are used to model the fan wake. Figure 17 shows the instantaneous entropy reconstructed around the $75\%$ span of the fan and it can be seen that the fan wake is transmitted into the computational domain of OGV without attenuation on the fan-OGV interface. Figure 17 also shows the instantaneous entropy distribution in an axial plane located close to the OGV inlet. It can be seen that the fan wake is correctly reconstructed in the computational domain of OGV and ESS. As the focus of the paper is on the computational performance of the generalized RK scheme, more detailed quantitative validations of the URANS will be presented in future work.
Table 5: FNLH setup for the low pressure compression system of a turbofan
Bladerow Temporal mode Spatial mode Note
Fan $-i\text{N}_{ESS}\Omega_{Fan},i\in\{1,2,3,4\}$ $i\text{N}_{ESS},i\in\{1,2,3,4\}$ exit, ESS potential field
$-i\text{N}_{ESS}\Omega_{Fan},i\in\{1,2,3,4\}$ $i\text{N}_{OGV},i\in\{1,2,3,4\}$ exit, OGV potential field
ESS $i\text{N}_{Fan}\Omega_{Fan},i\in\{1,2,...8\}$ $i\text{N}_{Fan},i\in\{1,2,...8\}$ inlet, Fan wake
OGV $i\text{N}_{Fan}\Omega_{Fan},i\in\{1,2,...8\}$ $i\text{N}_{Fan},i\in\{1,2,...8\}$ inlet, Fan wake
Figure 17: Reconstructed instantaneous entropy of the turbo-fan LPC system.
Figure 18 shows the comparisons of the convergence history of explicit and implicit RK schemes. 40 CPU cores are used. The CFL number of the implicit scheme is 50 and 4 levels of MG and V-cycle are used by the explicit scheme. It can be seen that the implicit scheme shows roughly 8x speed-up compared to the explicit scheme in terms of computational time. This is consistent with the previous two cases. In addition, the residuals of the explicit schemes stagnate to a higher level, while the implicit scheme shows a better convergence rate and can reach a deeper convergence compared to the explicit scheme. This shows that for challenging industrial cases, the implicit scheme not only shows a faster convergence rate, but also reaches a deeper convergence compared to the explicit scheme.
Figure 18: Comparisons of explicit and implicit schemes for the LPC system.
@@
Table 6 compares the memory consumption of the explicit and implicit RK schemes for the test cases. It can be seen that the memory consumption of the implicit solver is roughly half of the explicit solver using 4 MG levels. Therefore, the implicit RK scheme not only delivers faster and deeper convergence compared to the explicit scheme, but also has significantly less memory consumption. The memory consumption of the corresponding whole annulus URANS simulations is also summarized in Table 6. The memory consumption of FNLH is roughly two orders of magnitude smaller compared to the whole annulus URANS.
@@
Case 1 3.9 2.1 309.6
Case 2 7.4 3.2 378.1
Case 3 56.5 32.6 960.2
Figure 19 shows the convergence history of the implicit scheme (CFL = 50) for the test cases, but the x-axis of the plots is the normalized time, which is defined as the wall clock time of the FNLH simulation divided by the wall clock time of the execution of the full annulus URANS simulation for one revolution. It is typical to run 4-5 revolutions before collecting data from URANS. Figure 19 shows that the implicit FNLH is at least two orders of magnitude more efficient than the corresponding full-annulus URANS simulations in all test cases.
@@
Conclusions and Future Work
A unified formulation has been developed and implemented in the industrial CFD solver Hydra to accelerate the convergence of the FNLH method. A comparative study of the resulting explicit and implicit schemes has been undertaken. The proposed method is memory efficient as memory consumption depends weakly on the number of harmonics to compute. For the implicit scheme, the complex-number representation of the matrix is used for the matrix factorization for the purpose of reusing the existing lookup table for that of the mean flow. This significantly reduces the effort of implementing the implicit FNLH scheme.
The implicit FNLH scheme is found to consistently show a speedup of 7-10 compared to the explicit schemes with multigrid and this speed-up is comparable to the value reported by Misev [25] for the baseline steady flow solver. This demonstrates the optimal performance of the current implementation of the implicit FNLH method. Furthermore, the memory consumption of the implicit scheme consumes approximately $50\%$ of that of the explicit scheme with 4 levels of multigrid. The parallel performance of the implicit scheme is also studied and shows that the implicit solver is slightly less efficient as more CPU cores are used.
Future work includes removing the restriction on frozen turbulence and considering the cross-coupling of the harmonics. These treatments could have an impact on the convergence and robustness of the FNLH method. Besides, the robustness of the proposed procedure for cases with strong flow separation is not fully investigated in this paper. As the flow approaches the stall condition, it can generate unsteady flow disturbances whose frequencies are not relevant to the shaft speed. As harmonic methods require the specification of temporal and spatial modes a-priori, this poses further challenges to harmonic methods. Therefore, the implicit solver needs to be enhanced to overcome this challenge, and this will be investigated in future work.